Judy invests a certain amount of money at 7% annual interest and three times that amount at 8%. If her annual income is $232.50, how much does she have invested at each rate?
Thank you for your time!

Answer by checkley71(8403)

.07X+.08*3X=232.50
.07X+.24X=232.50
.31X=232.50
X=232.50/.31
X=750
3X=2250
PROOF
.07*750+.08*2250=232.50
52.50+180=232.50
232.50=232.50